what is the period of |sin3x|+|cos3x|

Grade:12

2 Answers

anvesh
8 years ago
  1. period of mod sinx is ^ and same for cosx so period is period of sin3x and cos3x lcm so answer is lcm of ^/3 and ^/3 that is ^/3
  2.  
Akshay
8 years ago
@anvesh: Not necessary, Assume a function f(x) = |sin x| + |cos x|
it will be repeated after pi/2, as | cos(pi/2+x) | = | sin x | and same for | sin(pi/2+x) | = | cos x |. 
So, f(x+pi/2) = f(x).
So, period here is pi/6 and not pi/3.